solve : 3x+5y=7 and 4x-3y=-10. if (a,b) is the solution, what is the sum of a and b?
Answer by stanbon(75887)

solve : 3x+5y=7 and 4x-3y=-10. if (a,b) is the solution, what is the sum of a and b?
--------------------
3x + 5y = 7
4x - 3y = -10
-------
Modify for elination:
12x + 20y = 28
12x - 9y = -30
-----------------------
Subtract and solve for "y":
29y = 58
y = 2
-----
Solve for "x":
3x + 5y = 7
3x + 10 = 7
3x = -3
-----
x = -1
------------
(a,b) = (-1,2)
----
Ans: a*b = -2
====================
Cheers,
Stan H.